# Parsimonious Compactly Supported Covariance Models in the Gauss Hypergeometric Class: Identifiability, Reparameterizations, and Asymptotic Properties

### v2

@@ A.5 Proof of Theorem 5
Let $\delta=\frac{d}{2}+\kappa$. Since $\log$ is strictly increasing, maximizing $A_{\kappa,\mu,l,a,d}$ is equivalent to maximizing
$LA_{\kappa,d}(\mu,l):=\log A_{\kappa,\mu,l,d,1},$
@@
By Theorem 4, for each fixed $l$ the maximizer in $\mu$ is $\mu=\mu_{1}(l)$ for $0\leq l\leq\delta$ and $\mu=\mu_{2}(l)$ for $l>\delta$. Hence it suffices to study the two one–variable functions
$F_{1}(l):=LA_{\kappa,d}(\mu_{1}(l),l)\quad\text{on }[0,\delta],\qquad F_{2}(l):=LA_{\kappa,d}(\mu_{2}(l),l)\quad\text{on }(\delta,\infty).$
@@
Case $i=1$ ($0\leq l\leq\delta$).
For $0\leq l\leq\delta$, $\mu_{1}(l)=\delta-l+\tfrac{1}{2}$. Set
$u(l):=\frac{\mu_{1}(l)+1}{2}+\kappa=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}-\frac{l}{2}.$
Using the explicit form of $A_{\kappa,\mu,l,d,1}$, $F_{1}(l)$ differs from
$\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big)$
@@
$F_{1}^{\prime}(l)$ $=u^{\prime}(l)\Big[\psi(u(l))+\psi(u(l)+l)-\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big]$
$\qquad+\psi(u(l)+l)-\psi\Big(u(l)+l+\frac{d}{2}\Big),$
@@
$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(u(l)+l)-\psi(u(l))+\psi\Big(u(l)+\frac{d}{2}\Big)-\psi\Big(u(l)+l+\frac{d}{2}\Big)\Big].$
@@
$x:=\frac{d}{4}+\frac{3\kappa}{2}+\frac{3}{4}\quad\Longrightarrow\quad u(l)=x-\frac{l}{2},\qquad u(l)+l=x+\frac{l}{2}.$
@@
$F_{1}^{\prime}(l)=\frac{1}{2}\Big[\psi(x+\tfrac{l}{2})-\psi(x-\tfrac{l}{2})+\psi(x-\tfrac{l}{2}+\tfrac{d}{2})-\psi(x+\tfrac{l}{2}+\tfrac{d}{2})\Big]=\frac{1}{2}\Big(\Delta_{l/2}(x)-\Delta_{l/2}(x+\tfrac{d}{2})\Big).$
Since $\Delta_{l/2}$ is decreasing and $x+\tfrac{d}{2}>x$, we have $\Delta_{l/2}(x)\geq\Delta_{l/2}(x+\tfrac{d}{2})$, hence $F_{1}^{\prime}(l)\geq 0$ on $0,\delta]$. Thus $F_{1}$ is non-decreasing on $0,\delta]$ and its maximum is attained at $l=\delta$.
Case $i=2$ ($l>\delta$).
For $l>\delta$, $\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l$. Define
@@
Then $u(l)=x(l)-\tfrac{l}{2}$ and $u(l)+l=x(l)+\tfrac{l}{2}$. As above, up to an additive constant independent of $l$,
$F_{2}(l)=\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\Big(u(l)+\frac{d}{2}\Big)-\log\Gamma\Big(u(l)+l+\frac{d}{2}\Big).$
Differentiating (same computation as in Case $i=1$ but with general $u$) gives the identity
$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big)-\frac{1}{2}\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big),$
i.e.
$F_{2}^{\prime}(l)=\Big(x^{\prime}(l)-\frac{1}{2}\Big)\Big(\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\Big).$
Now observe that
$x^{\prime}(l)=\frac{1}{2}\cdot\frac{l}{\sqrt{l^{2}+(2\kappa+d+1)}}\in\Big(0,\frac{1}{2}\Big),$
so $x^{\prime}(l)-\frac{1}{2}<0$ for all $l>\delta$. On the other hand, as in Case $i=1$, concavity of $\psi$ implies $\Delta_{l/2}(x(l))-\Delta_{l/2}(x(l)+\tfrac{d}{2})\geq 0$. Therefore $F_{2}^{\prime}(l)\leq 0$ for all $l>\delta$. Hence $F_{2}$ is non-increasing on $(\delta,\infty)$ and
$\sup_{l>\delta}LA_{\kappa,d}(\mu_{2}(l),l)=\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l),$
equivalently
$\sup_{l>\delta}A_{\kappa,\mu_{2}(l),l,a,d}=\lim_{l\downarrow\delta}A_{\kappa,\mu_{2}(l),l,a,d}.$
Finally, $\mu_{2}(l)\to 1$ as $l\downarrow\delta$ and $\mu_{1}(\delta)=1$. By continuity of $LA_{\kappa,d}(\mu,l)$ on the admissible region,
$\lim_{l\downarrow\delta}LA_{\kappa,d}(\mu_{2}(l),l)=LA_{\kappa,d}(1,\delta)=LA_{\kappa,d}(\mu_{1}(\delta),\delta).$
Therefore the maximum of $A_{\kappa,\mu_{1}(l),l,a,d}$ over $0\leq l\leq\delta$ is attained at $l=\delta$, and the supremum of $A_{\kappa,\mu_{2}(l),l,a,d}$ over $l>\delta$ equals the same value and is approached as $l\downarrow\delta$. Consequently, the global maximum over the admissible ranges is achieved at $(\mu,l)=(1,\delta)$, while in the regime $l>\delta$ it is not attained but only approached in the limit $l\downarrow\delta$. This proves Theorem 5. ∎
@@ A.6 Proof of Theorem 6
Using (7) and Equation (25) in Emery and Alegría (2022), the Fourier transform of the right-hand side member of (18) is found to be
$\begin{split}&\frac{\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)\Gamma(\delta)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{1}{2})\Gamma(2\delta)\Gamma(\frac{\mu-1}{2})\Gamma(\frac{\mu-1-d}{2}-\kappa+l)}\\ &\times\int_{0}^{1}\int_{0}^{1}u^{\delta-\frac{1}{2}}(1-u)^{\frac{\mu-3}{2}}v^{d+2\kappa}(1-v)^{\frac{\mu-d-3}{2}-\kappa+l}\mathstrut_{1}F_{2}\Big(\delta;\delta+\frac{1}{2},2\delta;-\frac{uv(za)^{2}}{4}\Big){\rm d}u{\rm d}v\\ &=\frac{\Gamma(\delta)\Gamma(\delta+\frac{\mu-d}{2})\Gamma(\delta+\frac{\mu-d}{2}+l)a^{d}}{2^{d}\pi^{\frac{d}{2}}\Gamma\left(\delta-\frac{d}{2}\right)\Gamma(\delta+\frac{\mu}{2})\Gamma(\delta+\frac{\mu}{2}+l)}\mathstrut_{1}F_{2}\Big(\delta;\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l;-\frac{(za)^{2}}{4}\Big)\\ &=\widehat{\mathcal{GH}}_{\delta,\delta+\frac{\mu}{2},\delta+\frac{\mu}{2}+l,a}(z),\quad z\geq 0.\end{split}$
The claim of the theorem is obtained by inverse Fourier transformation. ∎
@@ A.9 Proof of Theorem 9
Let $\lambda=\frac{1}{2}+\delta-\beta-\gamma.$ Note that under parameterization (33),
$\lambda=\frac{1}{2}+\delta-\beta-\gamma=\frac{1}{2}-\delta-\frac{d}{2}-\kappa=-\mu-\frac{d}{2}-\kappa-\frac{1}{2}<0,$
and also $\beta+\gamma=3\delta+\mu/2+d/2+\kappa-1/2$, so that $\beta+\gamma>(d+1)/2+3\delta$ holds whenever $\mu\geq 1$, which is exactly the validity condition of the $\mathcal{H}$ model. Hence Lemma 2 applies.
@@
$\qquad\times\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}-1\Bigg|,$
@@
$T_{3}(z)=w_{2}z^{\lambda}\{\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}.$
@@
For $T_{3}$, since $\lambda<0$, the dominant contribution of $T_{3}$ comes from the $z^{2\nu+d}$ factor:
$T_{3}(z)=w_{2}\,z^{\lambda+2\nu+d}\{\cos(\beta z-c_{5})+\mathcal{O}(z^{-1})\}+\mathcal{O}(z^{\lambda+2\nu+d-2}).$
Under $2\nu+d=2\delta$, we have $\lambda+2\nu+d=\lambda+2\delta$. Substituting $\lambda=-\mu-d/2-\kappa-1/2$ and $2\delta=d+1+2\kappa$,
$\lambda+2\delta=\frac{1}{2}-\mu-\frac{d}{2}+\kappa<0,$
since $\mu\geq 1$ and $\kappa>-1/2$. Hence $T_{3}(z)=\mathcal{O}(z^{\lambda+2\delta})$ with $\lambda+2\delta<0$, and by the Riemann–Lebesgue argument for oscillatory integrals,
$\int_{c}^{\infty}z^{d-1}T_{3}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-1+2(\lambda+2\delta)}\,\mathrm{d}z\right),$
which converges provided $d-1+2(\lambda+2\delta)<-1$, i.e. $d+2\lambda+4\delta<0$. Substituting,
$d+2\lambda+4\delta=d+2\!\left(-\mu-\tfrac{d}{2}-\kappa-\tfrac{1}{2}\right)+2(d+1+2\kappa)=d+1+2\kappa-2\mu.$
Since $\mu\geq 1$ and $\kappa>-1/2$, this quantity is at most $d-1+2\kappa<d$. For $d=1,2,3$ and $\mu\geq 1$, one verifies directly that $d+1+2\kappa-2\mu\leq d-1+2\kappa<d$; finiteness holds strictly when $\mu>(d+2\kappa)/2$, which is implied by the validity condition $\mu\geq 1$ for $d=1,2,3$ and $\kappa>-1/2$.
@@
Combining the above, condition (34) holds whenever $\nu=\kappa+\frac{1}{2}$, $w_{1}=1$, and $d=1,2,3$. Hence
@@
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}B(\kappa,\mu,d),$
where $B(\kappa,\mu,d)$ is defined in (20). Finally, the necessary part follows from Zhang (2004), by the same argument as in the proof of Theorem 1. ∎
A.10 Proof of Lemma 1

### v3

@@ A.5 Proof of Theorem 5
Let $\tau=\frac{d}{2}+\kappa$. Since $\log$ is strictly increasing, maximizing $A_{\kappa,\mu,l,a,d}$ is equivalent to maximizing
$LA_{\kappa,d}(\mu,l):=\log A_{\kappa,\mu,l,1,d},$
@@
By Theorem 4, for each fixed $l$ the maximizer in $\mu$ is $\mu=\mu_{1}(l)$ for $0\leq l\leq\tau$ and $\mu=\mu_{2}(l)$ for $l>\tau$, where
$\mu_{1}(l)=\tau+1-l\qquad(0\leq l\leq\tau),\qquad\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l\qquad(l>\tau).$
Hence it suffices to study the two one-variable functions
$F_{1}(l):=LA_{\kappa,d}(\mu_{1}(l),l)\quad\text{on }[0,\tau],\qquad F_{2}(l):=LA_{\kappa,d}(\mu_{2}(l),l)\quad\text{on }(\tau,\infty).$
@@
Case $i=1$ ($0\leq l\leq\tau$).
For $0\leq l\leq\tau$, $\mu_{1}(l)=\tau+1-l$. Set
$u(l):=\frac{\mu_{1}(l)+1}{2}+\kappa=\frac{d}{4}+\frac{3\kappa}{2}+1-\frac{l}{2}.$
Using the explicit form of $A_{\kappa,\mu,l,1,d}$, $F_{1}(l)$ differs from
$\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\!\left(u(l)+\frac{d}{2}\right)-\log\Gamma\!\left(u(l)+l+\frac{d}{2}\right)$
@@
$F_{1}^{\prime}(l)$ $=u^{\prime}(l)\left[\psi(u(l))+\psi(u(l)+l)-\psi\!\left(u(l)+\frac{d}{2}\right)-\psi\!\left(u(l)+l+\frac{d}{2}\right)\right]$
$\qquad+\psi(u(l)+l)-\psi\!\left(u(l)+l+\frac{d}{2}\right),$
@@
$F_{1}^{\prime}(l)=\frac{1}{2}\left[\psi(u(l)+l)-\psi(u(l))+\psi\!\left(u(l)+\frac{d}{2}\right)-\psi\!\left(u(l)+l+\frac{d}{2}\right)\right].$
@@
$x:=\frac{d}{4}+\frac{3\kappa}{2}+1\quad\Longrightarrow\quad u(l)=x-\frac{l}{2},\qquad u(l)+l=x+\frac{l}{2}.$
@@
$F_{1}^{\prime}(l)=\frac{1}{2}\!\left[\psi\!\left(x+\tfrac{l}{2}\right)-\psi\!\left(x-\tfrac{l}{2}\right)+\psi\!\left(x-\tfrac{l}{2}+\tfrac{d}{2}\right)-\psi\!\left(x+\tfrac{l}{2}+\tfrac{d}{2}\right)\right]=\frac{1}{2}\!\left(\Delta_{l/2}(x)-\Delta_{l/2}\!\left(x+\tfrac{d}{2}\right)\right).$
Since $\Delta_{l/2}$ is decreasing and $x+\tfrac{d}{2}>x$, we have $\Delta_{l/2}(x)\geq\Delta_{l/2}(x+\tfrac{d}{2})$, hence $F_{1}^{\prime}(l)\geq 0$ on $0,\tau]$. Thus $F_{1}$ is non-decreasing on $0,\tau]$ and its maximum is attained at $l=\tau$.
Case $i=2$ ($l>\tau$).
For $l>\tau$, $\mu_{2}(l)=\sqrt{l^{2}+(2\kappa+d+1)}-l$. Define
@@
Then $u(l)=x(l)-\tfrac{l}{2}$, $u(l)+l=x(l)+\tfrac{l}{2}$, and
$x^{\prime}(l)=\frac{1}{2}\cdot\frac{l}{\sqrt{l^{2}+(2\kappa+d+1)}}\in\left(0,\frac{1}{2}\right),\qquad u^{\prime}(l)=x^{\prime}(l)-\frac{1}{2},\qquad\frac{d}{dl}(u(l)+l)=x^{\prime}(l)+\frac{1}{2}.$
As above, up to an additive constant independent of $l$,
$F_{2}(l)=\log\Gamma(u(l))+\log\Gamma(u(l)+l)-\log\Gamma\!\left(u(l)+\frac{d}{2}\right)-\log\Gamma\!\left(u(l)+l+\frac{d}{2}\right).$
A direct differentiation, using $u^{\prime}(l)=x^{\prime}(l)-\tfrac{1}{2}$ and $\tfrac{d}{dl}(u(l)+l)=x^{\prime}(l)+\tfrac{1}{2}$, gives
$F_{2}^{\prime}(l)=\Bigl(x^{\prime}(l)-\tfrac{1}{2}\Bigr)\bigl[\psi(u)-\psi\!\left(u+\tfrac{d}{2}\right)\bigr]+\Bigl(x^{\prime}(l)+\tfrac{1}{2}\Bigr)\bigl[\psi(u+l)-\psi\!\left(u+l+\tfrac{d}{2}\right)\bigr],$ (34)
where $u=u(l)$. To determine the sign of $F_{2}^{\prime}(l)$, set
$h(t):=\psi(t)-\psi\!\left(t+\frac{d}{2}\right),\qquad t>0.$
Since $\psi$ is concave, $\psi^{\prime}$ is decreasing, hence $h^{\prime}(t)=\psi^{\prime}(t)-\psi^{\prime}(t+\tfrac{d}{2})\geq 0$, so $h$ is non-decreasing. Moreover $h(t)<0$ for all $t>0$ since $\psi$ is increasing. Because $u<u+l$, it follows that
$h(u)\leq h(u+l)<0.$
Since $x^{\prime}(l)-\tfrac{1}{2}<0$, replacing $h(u)$ by the larger value $h(u+l)$ in the first term of (34) gives the upper bound
$F_{2}^{\prime}(l)\leq\Bigl(x^{\prime}(l)-\tfrac{1}{2}\Bigr)h(u+l)+\Bigl(x^{\prime}(l)+\tfrac{1}{2}\Bigr)h(u+l)=2x^{\prime}(l)\,h(u+l)\leq 0.$
Therefore $F_{2}$ is non-increasing on $(\tau,\infty)$ and $\sup_{l>\tau}A_{\kappa,\mu_{2}(l),l,a,d}=\lim_{l\downarrow\tau}A_{\kappa,\mu_{2}(l),l,a,d}.$ Since $\mu_{2}(l)\to 1$ as $l\downarrow\tau$ and $\mu_{1}(\tau)=1$, by continuity of $LA_{\kappa,d}(\mu,l)$ on the admissible region,
$\lim_{l\downarrow\tau}LA_{\kappa,d}(\mu_{2}(l),l)=LA_{\kappa,d}(1,\tau)=LA_{\kappa,d}(\mu_{1}(\tau),\tau).$
Therefore the maximum of $A_{\kappa,\mu_{1}(l),l,a,d}$ over $0\leq l\leq\tau$ is attained at $l=\tau$, and the supremum of $A_{\kappa,\mu_{2}(l),l,a,d}$ over $l>\tau$ equals the same value and is approached as $l\downarrow\tau$. Consequently, the global maximum over the admissible region is achieved at $(\mu,l)=(1,\tau)=(1,d/2+\kappa)$, while in the regime $l>\tau$ it is not attained but only approached in the limit $l\downarrow\tau$. Restoring the factor $a^{d}$ yields the statement. ∎
@@ A.6 Proof of Theorem 6
Since all models are compactly supported and hence absolutely integrable, it suffices to verify the identity at the level of spectral densities. Taking the Fourier transform of both sides of (18) and using the scaling identity $\widehat{\mathcal{H}}_{\kappa,1,a\sqrt{uv},d}(z)=(a\sqrt{uv})^{d}\,\widehat{\mathcal{H}}_{\kappa,1,1,d}(za\sqrt{uv})$, the right-hand side of (18) has spectral density
$I(z)=C\,a^{d}\int_{0}^{1}\!\!\int_{0}^{1}u^{A}(1-u)^{B}\,v^{P}(1-v)^{Q}\,(uv)^{d/2}\,\widehat{\mathcal{H}}_{\kappa,1,1,d}(za\sqrt{uv})\,\mathrm{d}u\,\mathrm{d}v,$ (35)
where, from the exponents in (18),
$A=\delta-\tfrac{d+1}{2},\quad B=\tfrac{\mu-3}{2},\quad P=\delta-\tfrac{1}{2}+\kappa,\quad Q=\tfrac{\mu-d-3}{2}-\kappa+l.$
From (7), the spectral density of $\mathcal{H}_{\kappa,1,1,d}$ is
$\widehat{\mathcal{H}}_{\kappa,1,1,d}(w)=K\,{}_{1}F_{2}\!\left(\delta;\,\delta+\tfrac{1}{2},\,2\delta;\,-\tfrac{w^{2}}{4}\right)=K\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\delta+\frac{1}{2})_{k}\,(2\delta)_{k}}\frac{(-w^{2}/4)^{k}}{k!},$
where $K>0$ is a constant depending only on $(\kappa,d)$. Substituting $w=za\sqrt{uv}$ and inserting into (35):
$I(z)=C\,K\,a^{d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\delta+\frac{1}{2})_{k}\,(2\delta)_{k}}\frac{(-z^{2}a^{2}/4)^{k}}{k!}\;\mathcal{B}_{u}(k)\;\mathcal{B}_{v}(k),$
where the two integrals separate as
$\mathcal{B}_{u}(k):=\int_{0}^{1}u^{A+d/2+k}(1-u)^{B}\,\mathrm{d}u,\qquad\mathcal{B}_{v}(k):=\int_{0}^{1}v^{P+d/2+k}(1-v)^{Q}\,\mathrm{d}v.$
By the standard identity $\int_{0}^{1}t^{\alpha+k-1}(1-t)^{\beta-1}\,\mathrm{d}t=B(\alpha,\beta)\,(\alpha)_{k}/(\alpha+\beta)_{k}$, one obtains
$\mathcal{B}_{u}(k)$ $=B\!\left(A+\tfrac{d}{2}+1,\,B+1\right)\cdot\frac{(A+\frac{d}{2}+1)_{k}}{(A+B+\frac{d}{2}+2)_{k}},$ (36)
$\mathcal{B}_{v}(k)$ $=B\!\left(P+\tfrac{d}{2}+1,\,Q+1\right)\cdot\frac{(P+\frac{d}{2}+1)_{k}}{(P+Q+\frac{d}{2}+2)_{k}}.$ (37)
Substituting $A,B,P,Q$ and $\delta=\frac{d+1}{2}+\kappa$:
$A+\tfrac{d}{2}+1=\delta-\tfrac{d+1}{2}+\tfrac{d}{2}+1=\delta+\tfrac{1}{2},\quad A+B+\tfrac{d}{2}+2=\delta-\tfrac{d+1}{2}+\tfrac{\mu-3}{2}+\tfrac{d}{2}+2=\delta+\tfrac{\mu}{2}=\beta,$
$P+\tfrac{d}{2}+1=\delta-\tfrac{1}{2}+\kappa+\tfrac{d}{2}+1=2\delta,\quad P+Q+\tfrac{d}{2}+2=\delta-\tfrac{1}{2}+\kappa+\tfrac{\mu-d-3}{2}-\kappa+l+\tfrac{d}{2}+2=\delta+\tfrac{\mu}{2}+l=\gamma.$
Hence (36) and (37) become
$\mathcal{B}_{u}(k)=B\!\left(\delta+\tfrac{1}{2},\,\tfrac{\mu-1}{2}\right)\frac{(\delta+\frac{1}{2})_{k}}{(\beta)_{k}},\qquad\mathcal{B}_{v}(k)=B\!\left(2\delta,\,\gamma-2\delta\right)\frac{(2\delta)_{k}}{(\gamma)_{k}}.$
Inserting back into the power series, the factors $(\delta+\tfrac{1}{2})_{k}$ and $(2\delta)_{k}$ cancel with those in the denominator of the $\mathcal{H}_{\kappa,1}$ expansion, leaving
$I(z)=C^{\prime}\,a^{d}\sum_{k=0}^{\infty}\frac{(\delta)_{k}}{(\beta)_{k}\,(\gamma)_{k}}\frac{(-z^{2}a^{2}/4)^{k}}{k!}=C^{\prime}\,\widehat{\mathcal{GH}}_{\delta,\beta,\gamma,a}(z),$
where $C^{\prime}=C\cdot K\cdot B(\delta+\frac{1}{2},\frac{\mu-1}{2})\cdot B(2\delta,\gamma-2\delta)$. Setting $C$ to be the reciprocal of this product of Beta constants establishes the identity.
The integrability of the kernel (i.e., that the exponents $B=(\mu-3)/2>-1$ and $Q=(\mu-d-3)/2-\kappa+l>-1$) is guaranteed respectively by $\mu>1$ and $\mu+2l>1+d+2\kappa$, which are precisely the hypotheses of the theorem. ∎
@@ A.9 Proof of Theorem 9
Let $\lambda=\frac{1}{2}+\delta-\beta-\gamma.$ Under parameterization (41),
$\lambda$ $=\frac{1}{2}+\delta-\!\left(\delta+\frac{\mu}{2}\right)-\!\left(\delta+\frac{\mu}{2}+\frac{d}{2}+\kappa\right)$
$=\frac{1}{2}-\delta-\mu-\frac{d}{2}-\kappa$
$=-d-2\kappa-\mu\;<\;0,$
and also
$\beta+\gamma\;=\;2\delta+\mu+\frac{d}{2}+\kappa\;=\;3\delta+\mu-\frac{1}{2},$
so that $\beta+\gamma\geq 3\delta+\frac{1}{2}$ holds whenever $\mu\geq 1$, which is exactly the validity condition of the $\mathcal{H}$ model. Hence Lemma 1 applies.
@@
$\qquad\times\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}-1\Bigg|,$
@@
$T_{3}(z)=w_{2}z^{\lambda}\{\mathcal{O}(z^{2\nu+d-2})+z^{2\nu+d}\}\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}.$
@@
For $T_{3}$, since $\lambda<0$, the dominant contribution comes from the $z^{2\nu+d}$ factor:
$T_{3}(z)=w_{2}\,z^{\lambda+2\nu+d}\{\cos(az-c_{5})+\mathcal{O}(z^{-1})\}+\mathcal{O}(z^{\lambda+2\nu+d-2}).$
Under $2\nu+d=2\delta$,
$\lambda+2\nu+d\;=\;\lambda+2\delta\;=\;(-d-2\kappa-\mu)+(d+1+2\kappa)\;=\;1-\mu.$
Since $\mu\geq 1+\frac{d}{2}$ (as required by Theorems 10), we have $\lambda+2\delta=1-\mu<0$. By the Riemann–Lebesgue argument for oscillatory integrals,
$\int_{c}^{\infty}z^{d-1}T_{3}(z)^{2}\,\mathrm{d}z=\mathcal{O}\!\left(\int_{c}^{\infty}z^{d-1+2(1-\mu)}\,\mathrm{d}z\right),$
which converges provided $d-1+2(1-\mu)<-1$, i.e. $\mu>\frac{d}{2}+1$, which holds under the stated assumption.
@@
Combining the above, condition (42) holds whenever $\nu=\kappa+\frac{1}{2}$, $w_{1}=1$, $\mu\geq 1+\frac{d}{2}$, and $d=1,2,3$. Hence
@@
$\frac{\sigma_{0}^{2}}{\alpha^{2\nu}}=\frac{\sigma_{1}^{2}}{a^{2\kappa+1}}F(\kappa,\mu,d),$
where $F(\kappa,\mu,d)$ is defined in (20). Finally, the necessary part follows from Zhang (2004), by the same argument as in the proof of Theorem 1. ∎
A.10 Proof of Theorem 11
A preliminary lemma is required for the proof of Theorem 11. It establishes a uniform stochastic control of the plug-in microergodic estimator $\hat{\theta}_{n}(\mu,a)$ over $a\in I$, which is the key ingredient for showing that replacing a fixed $a$ by its ML estimator does not affect the fixed-domain limit distribution in Theorem 11.
Lemma 2.
Assume the setting of Theorem 10 with fixed $\mu>1+\frac{d}{2}$ and $a_{0}\in I$. Let
$\hat{\theta}_{n}(\mu,a)=\frac{\hat{\sigma}_{n}^{2}(\mu,a)\,F(\kappa,\mu,d)}{a^{2\kappa+1}},\qquad\hat{\sigma}_{n}^{2}(\mu,a)=\frac{\mathbf{Z}_{n}^{\top}R_{n}(\mu,a)^{-1}\mathbf{Z}_{n}}{n},$
and $\theta_{0}=\sigma_{0}^{2}F(\kappa,\mu,d)/a_{0}^{2\kappa+1}$. Then, as $n\to\infty$,
  1. (i)
    $\sup_{a\in I}\bigl|\hat{\theta}_{n}(\mu,a)-\theta_{0}\bigr|\xrightarrow{p}0$;
  2. (ii)
    for any fixed $a^{\star}\in I$, $\sup_{a\in I}\sqrt{n}\,\bigl|\hat{\theta}_{n}(\mu,a)-\hat{\theta}_{n}(\mu,a^{\star})\bigr|=o_{p}(1)$.